OpenStudy (anonymous):

Write the standard equation for the circle given the following equations x2 + y2 - 8x + 7 = 0

OpenStudy (anonymous):

x^2+y^2=r^2

OpenStudy (anonymous):

is that the formula?

OpenStudy (anonymous):

that is the general equation so now you need move terms around, this one would come out to x^2+y^2-8x=sqrt(-7)

OpenStudy (anonymous):

okay thanks could you help me with another problem similar to that one

OpenStudy (anonymous):

x2 + y2 + 4x - 6y - 3 = 0

OpenStudy (anonymous):

yeah just move the -3 over and then take square root of it

OpenStudy (anonymous):

now you dont have to, but to set it in the x^2+y^2=r^2 you can

OpenStudy (anonymous):

okay